OpenStudy (mathhelp346):

\[x ^{2}-2(3-x ^{2})\] Explain how you got your answer too :)

OpenStudy (mathhelp346):

??

OpenStudy (anonymous):

?

OpenStudy (mathhelp346):

what did they do?

OpenStudy (anonymous):

are you trying to simplify it?

OpenStudy (mathhelp346):

what were the steps they took to get 3x^2-6

OpenStudy (anonymous):

they factored it i believe

OpenStudy (mathhelp346):

how?

OpenStudy (anonymous):

kind of hard to explain

OpenStudy (jamesj):

@romoore: the fact that you can't explain it says everything we need to know about why using Wolframalpha is a bad substitute for thinking. x^2−2(3−x^2) = x^2 - 2(3) -2(-x^2) = x^2 - 6 + 2x^2 = 3x^2 - 6

OpenStudy (mathhelp346):

thanks!